OpenStudy (anonymous):

Find the vertex, focus, and directrix of the parabola and sketch its graph. Use a graphing utility to verify your graph. y^2-4x-12=0

OpenStudy (anonymous):

y^2-4x-12=0 given the form of the ecuation of the parabola (y-k)^2=4p(x-h) where (h,k) is the vertex and p is the distance from the vertice to the focus y^2=4x+12 ... Factoring 4 y^2=4(x+3) (y-0)^2=4(x-(-3)) here you can see: Vertex: (-3,0) Distance from the vertice to the focus and to the directriz: 4p=4 -> p=1 

OpenStudy (anonymous):

Thank you so much! Very informative.

OpenStudy (anonymous):

So the focus would be (-3,1) and the directrix would be y=3?

OpenStudy (anonymous):

directrix is y = -1

OpenStudy (anonymous):

sorry for the late

OpenStudy (anonymous):